Suppose the can segment their fans into young fans and senior citizens. Young fans have the demand curve P=120 - 10G (MR=120 - 20G). Senior citizens have the demand curve p=60 - 10G (MR=60 - 20G). Assume again that MC=0. What are the equilibrium price and quantity for young fans? What are the equilibrium price and quantity for senior citizens?
Answer #1

At equilibrium MR = MC

Young

120 - 20G = 0

G= 6

P= 120 - 10*6= 60

Equilibrium price and quantity for young are 60 and 6 units respectively.

Senior

60 - 20G = 0

G= 3

P= 60 - 10*3 = 30

Equilibrium price and quantity for seniors are 30 and 3 units respectively.